# Bridge Sampling Diagnostics

### v1

@@ 7 Discussion
We have proposed using the usual Monte Carlo standard error estimate to diagnose the reliability of (log) marginal likelihood estimates obtained by bridge sampling. We have demonstrated that this simple MCSE estimate is sufficiently accurate for practical purposes. When a MCSE estimate fails, it is likely that the provided MCSE estimate is anyway so big that it is indicating the need to run more MCMC iterations. In the case of high MCSE estimates, further accuracy can be obtained by running block reshuffling. The error in the log marginal likelihood seems to have a close to normal distribution, and the error in the marginal likelihood seems to have a close to log-normal distribution. Although the log-normal distribution has all finite moments, with finite sample size, the behavior can be indistinguishable from a distribution with infinite variance. The Pareto-$\hat{k}$ diagnostic can provide useful information about the reliability of MCSE estimates for bridge sampling, but its practical benefit is small due to the Pareto tail assumption not being sufficiently accurate for the relevant ratio distributions. We leave it for future research to investigate other tail shape measures that would work better with tail shapes arising in bridge sampling and could provide a more accurate diagnostic for how many more iterations would be needed for a desired accuracy of the (log) marginal likelihood estimate.
Acknowledgements
We thank Ozan Adıgüzel for making some of the initial experiments (not shown in this paper), and Noa Kallioinen and Anna Riha for helpful comments and feedback. We acknowledge the support by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence, Research Council of Finland project (340721), and The Aalto Science Institute international summer research internship programme, as well as the computational resources provided by the Aalto Science-IT project.
@@ Acknowledgements
We used 41 posteriors from posteriordb and a multi-component Gaussian process model (gpbf6), which was used by Zhang et al., (2022). Table A lists the posteriordb IDs, the posterior names used in the Figures in this paper, the number of posterior dimensions (in unconstrained space), and the short description of the model. The code for running the posterior inference and bridge sampling for all these and also for the simulated toy example is available at https://github.com/GiorgioMB/bridgesampling_paper_code.
@@
bball_drive_event_1-hmm_drive_1 Bball HMM 6 Hidden Markov model
bball_drive_event_0-hmm_drive_0 Bball0 HMM 6 Hidden Markov model
@@
eight_schools-eight_schools_centered 8-Schools CP 10 A centered hiearchical model for 8 schools
eight_schools-eight_schools_noncentered 8-Schools NCP 10 A non-centered hiearchical model for 8 schools
@@
fims_Aus_Jpn_irt-2pl_latent_reg_irt Japan IRT 531 Two-parameter logistic item response theory model
@@
radon_hierarchical_intercept_noncentered Radon Hierarchical 391 A county intercept and county level covariate for Radon data (non-centered)
@@
radon_variable_intercept_slope_noncentered Radon Intercept 777 Variable intercept and slope hiearchical for Radon data (noncentered)
@@
radon_partially_pooled_noncentered Radon Pooled 389 Hiearchical intercept model for Radon data (non-centered)
@@
radon_variable_slope_noncentered Radon Slope 390 Variable slope hiearchical for Radon data (non-centered)
@@
timssAusTwn_irt-gpcm_latent_reg_irt timssAusTwn IRT 530 Partial credit and generalized partial credit models with latent regression

### v2

@@ 7 Discussion
Our primary contribution is a diagnostic framework designed to quantify uncertainty at three distinct levels.
First, we derived and validated a closed-form Monte Carlo standard error (MCSE) for the bridge sampling estimator. While the asymptotic properties of bridge sampling are well-established (Meng and Wong,, 1996; Frühwirth-Schnatter,, 2004), practical reporting of uncertainty has been inconsistent. By verifying the MCSE against extensive empirical benchmarks from posteriordb, we demonstrate that for a range of posteriors, the standard delta-method approximation is sufficiently precise. This provides practitioners with a computationally cheap baseline diagnostic that should be standard reporting practice, akin to effective sample size (ESS) for MCMC.
Second, we bridged the gap between the theoretical boundedness of the estimator and its finite-sample behavior. While Meng and Wong, (1996) showed that the optimal bridge function yields bounded ratios, we identified a “pre-asymptotic regime” where these bounds are irrelevant for realistic sample sizes. By adapting the Pareto-$\widehat{k}$ diagnostic from importance sampling (Vehtari et al.,, 2024), we provide a method to detect when the estimator is dominated by rare events, extending the utility of tail-index diagnostics beyond unbounded importance sampling to the bounded (but heavy-tailed in behavior) setting of bridge sampling.
Third, we introduced block reshuffling as a bootstrapping-based diagnostic, allowing practitioners to assess the shape and skewness of the estimator’s sampling distribution without relying on the normality assumptions inherent in the standard analytic MCSE. While it may underestimate total variance due to sample reuse, it provides a critical check for non-normality and instability in finite samples.
We recommend a stepwise workflow: start with the analytic MCSE; if $\widehat{k}>0.5$ or if high precision is critical, verify stability with block reshuffling. This approach moves marginal likelihood estimation from a “black box” procedure to a transparent, diagnosable component of the Bayesian workflow.
8 Acknowledgments
We thank Ozan Adıgüzel for making some of the initial experiments (not shown in this paper), and Noa Kallioinen and Anna Riha for helpful comments and feedback. We also acknowledge the computational resources provided by the Aalto Science-IT project.
9 Funding
We acknowledge the support by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence, Research Council of Finland project (340721), and The Aalto Science Institute international summer research internship programme, as well as the computational resources provided by the Aalto Science-IT project.
10 Data Availability Statement
The data and code that support the findings of this study are available at the following URL: https://github.com/GiorgioMB/bridgesampling_paper_code. Monte Carlo standard error computation is also included in bridgesampling package https://github.com/quentingronau/bridgesampling/.
@@ 10 Data Availability Statement
We used 41 posteriors from posteriordb and a multi-component Gaussian process model (gpbf6), which was used by Zhang et al., (2022). Table A lists the posteriordb IDs, the posterior names used in the figures in this paper, the number of posterior dimensions (in unconstrained space), and a short description of the model.
Posteriors from posteriordb used in Section 6. Columns: posteriordb ID, name used in the plots, number of unconstrained posterior dimensions $D$ and a brief description.
@@
bball_drive_event_1-
hmm_drive_1
Bball HMM 6 Hidden Markov model
bball_drive_event_0-
hmm_drive_0
Bball0 HMM 6 Hidden Markov model
@@
eight_schools-eight_schools_
centered
8-Schools CP 10 A centered hierarchical model for 8 schools
eight_schools-eight_schools_noncentered 8-Schools NCP 10 A non-centered hierarchical model for 8 schools
@@
fims_Aus_Jpn_irt-2pl_latent
_reg_irt
Japan IRT 531 Two-parameter logistic item response theory model
@@
radon_hierarchical_intercept
_noncentered
Radon Hierarchical 391 A county intercept and county level covariate for Radon data (non-centered)
@@
radon_variable_intercept_
_slope_noncentered
Radon Intercept 777 Variable intercept and slope hierarchical for Radon data (noncentered)
@@
radon_partially_pooled_
_noncentered
Radon Pooled 389 Hierarchical intercept model for Radon data (non-centered)
@@
radon_variable_slope_
_noncentered
Radon Slope 390 Variable slope hierarchical for Radon data (non-centered)
@@
timssAusTwn_irt-gpcm_latent
_reg_irt
timssAusTwn IRT 530 Partial credit and generalized partial credit models with latent regression